# Debiasing astro-Photometric Observations with Corrections Using Statistics (DePhOCUS)

### v1

@@ 3.1.3 Observatory bias results
In the ATLAS c- and o-band the spread among the observatories is again quite low as expected and within $3\sigma$ of the standard error. There is only a slight deviation in c-band of around $0.03$ mag of T05 in Gaia-DR2 (V) compared to the other observatories. Also in o-band there is a similar deviation, but for T08.
For the Sloan filters g, r and i we can see that there is mostly only one major observatory per band and catalog, so the conclusions from Sec. 3.1.2 also apply to the aBCO results. As already mentioned, the significant deviation in r-band is caused by the deviation (and possible calibration) of G45. In i-band there is a slight effect of a $0.03$ mag difference for the Pan-STARRS telescopes F51 and F52. This difference can also be seen in the Pan-STARRS w-band. Additionally, there are some T08 and C41 measurements in the w-band, which deviate from the other observatories.
@@ 3.2 Correction systems
With the algorithm defined in Sec. 2.3.2 we generate different correction systems based on the bias results with different levels of probability $p$. These Debiased astro-Photometic Observations with Corrections Using Statistics (DePhOCUS) corrections are then compared to the results of previous corrections that are used.
The bias analysis results contain 20 band (B), 159 band-catalog (BC) and 2287 band-catalog-observatory (BCO) combinations, which we use as correction values. In the statistical correction algorithm, the standardized t-test is applied to all these combinations with $p$-levels of 00 (all), $0.68$ ($1\sigma$), $0.90$, $0.95$ ($2\sigma$), $0.99$ and $0.997$ ($3\sigma$), reducing the number of corrections with increasing significance. This approach keeps the highly statistical relevant corrections for high significance levels. In Tab. 3 we can see the decrease of correction rules from nearly 2500 to roughly 150 with increasing $p$.
For the 74 remaining out of the 468 asteroids from the asteroid photometry database [^10] https://www.asu.cas.cz/~ppravec/newres.htm by Pravec et al. (2012) that were not used for obtaining the biases, we can compute the RMS of the reduced magnitudes of all observations for each asteroid before and after applying the correction. Again, in order to avoid any effect of specific asteroids with comparatively many observations, the RMS is calculated for each asteroid and we look at the average RMS among all asteroids. The results for the average RMS for the different correction system (including existing corrections by MPC, NEOCC and JPL) are shown in Tab. 3.
Table 3: Summary of the results of the different DePhOCUS correction systems with $p$-level for the number of correction rules in band (B), band-catalog (BC) and band-catalog-observatory (BCO) combinations, and their effect on the RMS of observations from 74 independent validation objects with their relative change in RMS (assuming boundary of $0.25$ mag, cf. Sec. 2.3.3).
@@
$p=0$20 159 2287 0.3720 -30.3%
$p=0.68$17 90 701 0.3621 -35.9%
$p=0.90$17 61 339 0.3613 -36.4%
$p=0.95$16 53 241 0.3622 -35.9%
$p=0.99$12 41 128 0.3691 -32.0%
$p=0.997$10 34 91 0.3694 -31.8%
@@
For the NEOCC correction that was used prior to 2023-09-28 there are also similar issues, but the different correction in the i- and especially w-band reduce uncertainties, compensating the worsening and leading to an almost zero change compared to no correction. On the other hand, the current NEOCC correction, in use since 2023-09-28, includes the aB bias analysis results of the present study. Band corrections that fulfill the $p=0.9$ criterion are allowed to replace previous correction (cf. Tab. 4), which happens to be the case for 17 out of the 24 different bands. This correction, that is just using band correction, already reduces the overall RMS to $0.3768$ mag, a decrease of $27.6\%$ compared to $0.25$ mag estimated correction limit (cf. Sec. 2.3.3). The JPL correction, using the same band correction as the prior 2023-09-28 NEOCC-version, except for G-band, is also reducing the RMS by $4.0\%$, showing the impact of minor correction changes for commonly-used bands.
Table 4: Current NEOCC correction values (used since 2023-09-28), only using aB bias analysis results in correction algorithm ($p=0.9$). The results for the values with two decimal places do not fulfill the significance criteria, therefore the previous values are being used there.
@@
Introducing the full DePhOCUS correction, that is combining all the band, catalog and observatory bias results, we can see that the RMS is reducing by more than $30\%$ for all significance levels, having an optimum RMS of $0.3613$ mag ($-36.4\%$) for a significance level $p=0.90$. This already shows the large potential of catalog and observatory correction in addition to band correction. Using more or even all corrections (smaller $p$) increases the uncertainties, as corrections might not be statistically solid enough. Using less corrections (higher $p$) selects the more statistically relevant corrections, but also removes corrections that are improving the results. A significance level of $p=0.90$ is therefore our empirical optimum for the correction algorithm. This set of corrections can be found in the appendix.
@@ 3.3 Validation and prediction of correction model
We use the 74 reference asteroids to validate the results for the developed correction system and see if this leads to a better prediction for the $H$-$G$ magnitude system. We achieve this by blindly fitting the corrected astro-photometric observations to the $H$-$G$ phase curve and comparing the fitted parameters to the reference $H$-$G$ parameters by Pravec et al. (2012). In order to avoid specific biases from the fitting of the $G$ parameter, which is not well understood and rarely done up to this point, we constrain the $G$ parameter in the fit to the reference value $G_{\text{Pravec}}$ and do only optimize the phase curve function for $H_{\text{fit}}$. The fitting procedure is using a non-linear least squares method with Eq. (3) as fitting function, weighting the observations by their photometric precision [^11] As used by NEOCC: $\sigma=(0.5,0.7,1.0)$ mag for 2, 1 and 0 decimal places..
Fig. 9 shows the results for the fitted $H_{\text{fit}}$ and the resulting $RMS_{\text{fit}}$ in comparison to the reference $H_{\text{Pravec}}$, where we have a deviation $RMS_{\text{Pravec}}$ in the observation with the applied DePhOCUS correction ($p=0.9$). The results show that the $H_{\text{fit}}$ predicts well the reference values within a standard deviation of $0.106$ mag, having no significant systematic deviation or bias ($0.003$ mag), which shows there is no selection bias [^12] Could have been due to the selection by time (after 2021-10-21). of the reference asteroids. The results also show high agreement for the correlation of fitted and reference $H$ parameter ($R^{2}=0.997$) with a reduced $\chi^{2}$ value of $0.081$ for the fitting uncertainties and $2.461$ for the reference error estimates. We are aware that the fitted $H$ parameter uncertainties are underestimating the true uncertainties. As they are based on the RMS of the observation, the small fitting errors do not consider the intrinsic fluctuations of the parameter (e.g., due to rotational effects). We see that the RMS is decreasing on average by $0.013$ mag with the fitted parameters, with minor exceptions, as the fitted parameters should minimize deviations, leading to an expected low correlation factor $R^{2}=0.867$.
@@
Figure 9: Histogram of difference between fitted $H$ parameter, after applying DePhOCUS correction ($p=0.9$), and reference values $H_{\text{Pravec}}$ by Pravec et al. (left) with their correlation (right), when keeping $G=G_{\text{Pravec}}$ for 74 asteroids as validation. Also comparing the $RMS$ of measurements by using the fitted parameters or the reference values.
@@
The results for the $H$ parameter prediction with the DePhOCUS correction are significantly reducing uncertainties compared to previous corrections; the deviation of the fitted $H$ parameter to the reference values with MPC correction is $0.136$ mag for the same objects (cf. Appendix), resulting in a $28\%$ reduction of uncertainties with the new method. While the coefficient of correlation is already nominally high for the MPC method ($R^{2}=0.995$), we especially see improvements in the reduced $\chi^{2}$ value with a $21\%$ and $94\%$ increase in goodness of fit for the fitting uncertainties and the reference error estimates respectively. [^13] For MPC correction: $\chi^{2}_{f}=0.098$ and $\chi^{2}_{t}=4.770$.
@@
Figure 10: RMS distribution for 74 validation asteroids with DePhOCUS correction ($p=0.9$) and MPC correction.
@@
Figure 11: Distribution of difference in optimized $H$ for 74 validation asteroids with DePhOCUS correction ($p=0.9$) and MPC correction.
@@
Looking at Fig. 10, we can also compare the distribution of RMS for the 74 asteroid phase curves with the novel correction method in contrast to the current MPC correction (cf. Fig. 13). It is noticeable that the RMS distribution shifts to lower RMS values, on average by $-0.071$ mag, reaching a peak at around $0.35$ mag with the new method. The reduction is consistent at all levels of RMS. The distribution also gets narrower with the DePhOCUS method, especially reducing the number of high-RMS outliers. This leads to a reduction in the standard deviation from $\sigma_{\text{MPC}}=0.093$ down to $\sigma_{\text{DePhOCUS}}=0.080$.
@@ 4.1 Advanced bias analysis results
in comparison to the obtained correction values for the observatory stations. We achieve this by binning the data points of each station (typically $0.1$ mag for the bias) in a hexagonal structure and count the number of observations that fall in each hexagon. This visualizes the density of observations to be located at a certain parameter- and bias-area.
@@
The astrometric residuals of the observations are a metric for potential general mistakes in the measurement. This could be also due to trailing or an inaccurate measurement. In Fig. 19 we can see that many of the observatories have a tendency for an increasingly negative bias for higher astrometric residuals, especially codes 704 and F51, which match well with the trends from the apparent motion discussion and could be due to trailing. The results from codes T05, T08, G96 and C57 show no effect on the bias, even though the latter has a quite large range of astrometric errors compared to the others. Overall, it can be seen that code C57, despite having significant errors in the astrometry, does not seem to have systematic deviations in the astro-photometry141414apart from a small proportion of large outliers around $-3$ mag bias..
@@
The display of the evolution of the bias over time as shown in Fig. 22 is a measure to monitor long-term effects, but also abrupt changes. For codes 703 and G96 there is a significant step visible happening during the summer of 2018, where the bias drops by more than $0.3$/$0.1$ mag respectively. This drop perfectly corresponds to the switch from Gaia DR1 to DR2 as reference star catalog and matches well with the corresponding aBCO corrections obtained for the sites. This shows that effects that result from different catalogs being used are well covered by our algorithm. However, the increase in the bias for code G96 starting after the summer of 2022, even though still using Gaia DR2, is not included in the corrections. Also, more gradual effects like the steep increase of bias for code 704 in its first years of measurement up to the year 2000 and the consequent shallow decrease up to the year 2006 can be seen in the analysis. Also code F51 had high fluctuations in its first years of measurement, having a rather negative bias then, but afterwards it has, as well as code 704, a quite constant bias.151515with the exception of down-times during 2018 and 2022. For code T05 and T08 there is no change of bias visible over time.161616with the exception of a lack of measurements during 2022.
@@ 4.2 Additional points
It turns out that there are some highly significant bias results with much underlying data and high deviations to the previous correction. These have the largest individual impact on the results. On the other hand, there are many lower significant results with less data and only slight differences. These might be giving questionable results or not having large individual impact, but there could still be reasonable results among them, where we might miss their impact, when not including them. Finding the optimum point for the minimum RMS based on the different $p$-levels gives us the opportunity to include all relevant correction, but not to loose too many reasonable results at the same time.
@@ 5 Conclusion
In this paper we determined significant biases for the major color bands used for minor planet observations, showing that there is a large variation in accuracy in the field of astro-photometry. We were able to quantify the systematic deviations in different color band, reference star catalogs and for specific observatories and give reasons for some of the occurring issues. Therefore, we propose to use band-, catalog- and observatory specific bias corrections. Based on these bias results, we were able to develop a novel statistical correction algorithm and generated a correction system in order to debias all the astro-photometric observations. This led to a $51\%$ decrease in the RMS of the asteroids’ phase curve compared to current methods, shown for an independent set of objects. These were also used to validate our method for deriving the absolute magnitude $H$. With the newly corrected observations we were able to improve the goodness of fit statistics by $94\%$ and reduce the uncertainties for $H$ by $28\%$ and thus improving size estimates. There is only a small proportion of objects in the solar system with accurately measured photometry, therefore astro-photometry is the best guess for most objects, especially in time-critical domains, as planetary defense with imminent impactors, or large data models in Solar System Science. In total, the new method can be beneficial in estimating the physical properties and their accuracy of almost all minor planets.

### v2

@@ 3.1.3 Observatory bias results
In the ATLAS c- and o-band the spread among the observatories is again quite low as expected and within $3\sigma$ of the standard error. There is only a slight deviation in c-band of around $0.03$ mag of T05 in Gaia-DR2 (V) compared to the other observatories. Also in o-band there is a similar deviation, but also for M22 and especially W68 in Gaia-DR2. This shows slight discrepencies among the ATLAS telescopes.
For the Sloan filters g, r and i we can see that there is mostly only one major observatory per band and catalog, so the conclusions from Sec. 3.1.2 also apply to the aBCO results. As already mentioned, the significant deviation in r-band is caused by the deviation (and possible calibration) of G45. In i-band there is a slight effect of a $0.03$ mag difference for the Pan-STARRS telescopes F51 and F52 in Gaia-DR1 (U). This difference can also be seen in the Pan-STARRS w-band and for other catalogs (L and X), with F52 always having a higher bias than F51, but the amount of the discrepancy is varying among the bands and catalogs. Additionally, there are some T08 and C41 measurements in the w-band, which deviate from the other observatories.
@@ 3.2 Correction systems
With the algorithm defined in Sec. 2.3.2 we generate different correction systems based on the bias results with different levels of confidence $\gamma=1-\alpha$. These Debiased astro-Photometic Observations with Corrections Using Statistics (DePhOCUS) corrections are then compared to the results of previous corrections that are used.
The bias analysis results contain 20 band (B), 159 band-catalog (BC) and 2287 band-catalog-observatory (BCO) combinations, which we use as correction values. In the statistical correction algorithm, the standardized t-test is applied to all these combinations with confidence levels $\gamma$ of 00 (all), $0.68$ ($1\sigma$), $0.90$, $0.95$ ($2\sigma$), $0.99$ and $0.997$ ($3\sigma$), reducing the number of corrections with increasing confidence. This approach keeps the highly statistical relevant corrections for high confidence levels. In Tab. 3 we can see the decrease of correction rules from nearly 2500 to roughly 150 with increasing $\gamma$.
For the validation subset, consisting of 74 out of the 468 asteroids from the asteroid photometry database [^12] https://www.asu.cas.cz/~ppravec/newres.htm by Pravec et al. (2012) that were not used for obtaining the biases, we can compute the RMS of the reduced magnitudes of all observations for each asteroid before and after applying the correction. Again, in order to avoid any effect of specific asteroids with comparatively many observations, the RMS is calculated for each asteroid and we look at the average RMS among all asteroids. The results for the average RMS for the different correction system (including existing corrections by MPC, NEOCC and JPL) are shown in Tab. 3.
Table 3: Summary of the results of the different DePhOCUS correction systems with confidence level $\gamma$ for the number of correction rules in band (B), band-catalog (BC) and band-catalog-observatory (BCO) combinations, and their effect on the RMS of observations from 74 independent validation objects with their relative change in RMS (assuming boundary of $0.25$ mag, cf. Sec. 2.3.3).
@@
$\gamma=0$20 159 2287 0.3720 -30.3%
$\gamma=0.68$17 90 701 0.3621 -35.9%
$\gamma=0.90$17 61 339 0.3613 -36.4%
$\gamma=0.95$16 53 241 0.3622 -35.9%
$\gamma=0.99$12 41 128 0.3691 -32.0%
$\gamma=0.997$10 34 91 0.3694 -31.8%
@@
For the NEOCC correction that was used prior to 2023-09-28 there are also similar issues, but the different correction in the i- and especially w-band reduce uncertainties, compensating the worsening and leading to an almost zero change compared to no correction. On the other hand, the current NEOCC correction, in use since 2023-09-28, includes the aB bias analysis results of the present study. Band corrections that fulfill the $\gamma=0.9$ criterion are allowed to replace previous correction (cf. Tab. 4), which happens to be the case for 17 out of the 24 different bands. This correction, that is just using band correction, already reduces the overall RMS to $0.3768$ mag, a decrease of $27.6\%$ compared to $0.25$ mag estimated correction limit (cf. Sec. 2.3.3). The JPL correction, using the same band correction as the prior 2023-09-28 NEOCC-version, except for G-band, is also reducing the RMS by $4.0\%$, showing the impact of minor correction changes for commonly-used bands.
Table 4: Current NEOCC correction values (used since 2023-09-28), only using aB bias analysis results in correction algorithm ($\gamma=0.9$). The results for the values with two decimal places do not fulfill the significance criteria, therefore the previous values are being used there.
@@
Introducing the full DePhOCUS correction, that is combining all the band, catalog and observatory bias results, we can see that the RMS is reducing by more than $30\%$ for all confidence levels, having an optimum RMS of $0.3613$ mag ($-36.4\%$) for a confidence level $\gamma=0.90$. This already shows the large potential of catalog and observatory correction in addition to band correction. Using more or even all corrections (smaller $\gamma$) increases the uncertainties, as corrections might not be statistically solid enough. Using less corrections (higher $\gamma$) selects the more statistically relevant corrections, but also removes corrections that are improving the results. A confidence level of $\gamma=0.90$ is therefore our empirical optimum for the correction algorithm. This set of corrections can be found in the appendix.
@@ 3.3 Validation and prediction of correction model
Figure 9: Histogram of difference between fitted $H$ parameter, after applying DePhOCUS correction ($\gamma=0.9$), and reference values $H_{\text{Pravec}}$ by Pravec et al. (left) with their correlation (right), when keeping $G=G_{\text{Pravec}}$ for 74 asteroids as validation. Also comparing the $RMS$ of measurements by using the fitted parameters or the reference values.
We use the subset of 74 asteroids to validate the results for the developed correction system and see if this leads to a better prediction for the $H$-$G$ magnitude system. We achieve this by blindly fitting the corrected astro-photometric observations to the $H$-$G$ phase curve and comparing the fitted parameters to the reference $H$-$G$ parameters by Pravec et al. (2012). In order to avoid specific biases from the fitting of the $G$ parameter, which is not well understood and rarely done up to this point, we constrain the $G$ parameter in the fit to the reference value $G_{\text{Pravec}}$ and do only optimize the phase curve function for $H_{\text{fit}}$. The fitting procedure is using a non-linear least squares method with Eq. (3) as fitting function, weighting the observations by their photometric precision [^13] As used by NEOCC: $\sigma=(0.5,0.7,1.0)$ mag for 2, 1 and 0 decimal places..
Fig. 9 shows the results for the fitted $H_{\text{fit}}$ and the resulting $RMS_{\text{fit}}$ in comparison to the reference $H_{\text{Pravec}}$, where we have a deviation $RMS_{\text{Pravec}}$ in the observation with the applied DePhOCUS correction ($\gamma=0.9$). The results show that the $H_{\text{fit}}$ predicts well the reference values within a standard deviation of $0.106$ mag, having no significant systematic deviation or bias ($0.003$ mag), which shows there is no selection bias [^14] Could have been due to the selection by time (after 2021-10-21). of the reference asteroids. The results also show high agreement for the correlation of fitted and reference $H$ parameter ($R^{2}=0.997$) with a reduced $\chi^{2}$ value of $0.081$ for the fitting uncertainties and $2.461$ for the reference error estimates. We are aware that the fitted $H$ parameter uncertainties are underestimating the true uncertainties. As they are based on the RMS of the observation, the small fitting errors do not consider the intrinsic fluctuations of the parameter (e.g., due to rotational effects). We see that the RMS is decreasing on average by $0.013$ mag with the fitted parameters, with minor exceptions, as the fitted parameters should minimize deviations, leading to an expected low correlation factor $R^{2}=0.867$.
@@
The results for the $H$ parameter prediction with the DePhOCUS correction are significantly reducing uncertainties compared to previous corrections; the deviation of the fitted $H$ parameter to the reference values with MPC correction is $0.136$ mag for the same objects (cf. Appendix, Fig. 14), resulting in a $28\%$ reduction of uncertainties with the new method. While the coefficient of correlation is already nominally high for the MPC method ($R^{2}=0.995$), we especially see improvements in the reduced $\chi^{2}$ value with a $21\%$ and $94\%$ increase in goodness of fit for the fitting uncertainties and the reference error estimates respectively. [^15] For MPC correction: $\chi^{2}_{f}=0.098$ and $\chi^{2}_{t}=4.770$.
@@
Figure 10: RMS distribution for 74 validation asteroids with DePhOCUS correction ($\gamma=0.9$) and MPC correction.
@@
Figure 11: Distribution of difference in optimized $H$ for 74 validation asteroids with DePhOCUS correction ($\gamma=0.9$) and MPC correction.
@@
Looking at Fig. 10-11, we can also compare the distribution of RMS and difference of the $H$ parameter for the 74 asteroid phase curves with the novel correction method in contrast to the current MPC correction (cf. Fig. 14). The $H$ parameter changes in a range of $-0.10$ to $0.15$ mag (standard deviation $0.059$ mag) with the new method. On average, the $H$ values do not tend to change in one specific direction (average difference $0.001$ mag). But however, for the RMS it is noticeable that the distribution shifts to lower RMS values, on average by $-0.071$ mag, reaching a peak at around $0.35$ mag with the new method due to change and improvement of the individual $H$ parameters. The reduction is consistent at all levels of RMS. The distribution also gets narrower with the DePhOCUS method, especially reducing the number of high-RMS outliers. This leads to a reduction in the standard deviation from $\sigma_{\text{MPC}}=0.093$ down to $\sigma_{\text{DePhOCUS}}=0.080$.
@@ 4.1 Advanced bias analysis results
in comparison to the obtained correction values for the observatory stations. We achieve this by binning the data points of each station in a hexagonal grid (spacing of $0.1$ mag for the bias) and aggregate all observations that fall in each segment and compute its number of included points. This visualizes the density of observations to be located at a certain parameter- and bias-area.
@@
The astrometric residuals of the observations are a metric for potential general mistakes in the measurement. This could be also due to trailing or an inaccurate measurement. In Fig. 19 we can see that many of the observatories have a tendency for an increasingly negative bias for higher astrometric residuals, especially codes 704 and F51, which match well with the trends from the apparent motion discussion and could be due to trailing. The results from codes T05, T08, G96 and C57 show no effect on the bias, even though the latter has a quite large range of astrometric errors compared to the others. Overall, it can be seen that code C57, despite having significant errors in the astrometry, does not seem to have systematic deviations in the astro-photometry161616apart from a small proportion of large outliers around $-3$ mag bias..
@@
The display of the evolution of the bias over time as shown in Fig. 22 is a measure to monitor long-term effects, but also abrupt changes. For codes 703 and G96 there is a significant step visible happening during the summer of 2018, where the bias drops by more than $0.3$/$0.1$ mag respectively. This drop perfectly corresponds to the switch from Gaia DR1 to DR2 as reference star catalog and matches well with the corresponding aBCO corrections obtained for the sites. This shows that effects that result from different catalogs being used are well covered by our algorithm. However, the increase in the bias for code G96 starting after the summer of 2022, even though still using Gaia DR2, is not included in the corrections. Also, more gradual effects like the steep increase of bias for code 704 in its first years of measurement up to the year 2000 and the consequent shallow decrease up to the year 2006 can be seen in the analysis. Also code F51 had high fluctuations in its first years of measurement, having a rather negative bias then, but afterwards it has, as well as code 704, a quite constant bias.171717with the exception of down-times during 2018 and 2022. For code T05 and T08 there is no change of bias visible over time.181818with the exception of a lack of measurements during 2022.
@@ 4.2 Additional points
It turns out that there are some highly significant bias results with much underlying data and high deviations to the previous correction. These have the largest individual impact on the results. On the other hand, there are many lower significant results with less data and only slight differences. These might be giving questionable results or not having large individual impact, but there could still be reasonable results among them, where we might miss their impact, when not including them. Finding the optimum point for the minimum RMS based on the different confidence levels gives us the opportunity to include all relevant correction, but not to loose too many reasonable results at the same time.
@@ 5 Conclusion
In this paper we determined significant biases for the major color bands used for minor planet observations, showing that there is a large variation in accuracy in the field of astro-photometry. We were able to quantify the systematic deviations in different color band, reference star catalogs and for specific observatories and give reasons for some of the occurring issues. Therefore, we propose to use band-, catalog- and observatory specific bias corrections. Based on these bias results, we were able to develop a novel statistical correction algorithm and generated a correction system in order to debias all the astro-photometric observations. This led to a $36\%$ decrease in the RMS of the asteroids’ phase curve compared to current methods, shown for an independent set of objects. These were also used to validate our method for deriving the absolute magnitude $H$. With the newly corrected observations we were able to improve the goodness of fit statistics by $94\%$ and reduce the uncertainties for $H$ by $28\%$ and thus improving size estimates. These results can be further improved in future studies by including other more specific corrections and using a photometric weighting system based on the deviations of observatories obtained in this study. All in all, there is only a small proportion of objects in the solar system with accurately measured photometry, therefore astro-photometry is the best guess for most objects, especially in time-critical domains, as planetary defense with imminent impactors, or large data models in Solar System Science. Therefore, the new method can be beneficial in estimating the physical properties and their accuracy of almost all minor planets.